OpenStudy (anonymous):

Which property states that adding zero to a number does not change the number? Associative Property of Addition Identity Property of Addition Distribution Property Commutative Property of Addition

OpenStudy (anonymous):

Which expression shows the result of applying the Commutative Property of Multiplication to the expression 4 • 5 • (3 + 1)? 4 • 5 • (4) 5 • 4 • (3 + 1) 4 • (5 • 3 + 1) 4 • (5 • 3 + 5 • 1)

OpenStudy (anonymous):

Which expression will help make the following problem easier to solve using mental math? Arsha wants to buy 16 gallons of gasoline, which cost $2.10 per gallon. 16 • (2 + 0.1) 16(2) + 2(0.1) 2 • (16 + 0.1) 16(0.1) + 0.1(2)

OpenStudy (anonymous):

Identity refers to how to make something exactly like itself. I can do this by multiplying by one or ADDING zero.

OpenStudy (anonymous):

identity property a*e=a example: a=1 1*e=1 e=1-1 e=0

OpenStudy (anonymous):

When I have to drive a long distance to work, I am COMMUTing. Meaning that I move around. That's what it means in math too.

OpenStudy (unklerhaukus):

Identity Property \[a\star 0=a=0\star a\] Associative property\[a\star(b\star c)=(a\star b)\star c\] Distribution Property\[a\star(b+ c)=a\star b+ a\star c\] Commutative Property\[a\star b=b\star a\]